# Estimating the FDR of variable selection

### v2

@@ 1 Introduction
When deciding which variables to include in a statistical model, a data analyst commonly wants to balance multiple goals including predictive accuracy, accurate detection of signal variables, parsimony, and interpretability. Perhaps the best-known paradigm is to use a variable-selecting method such as the Lasso (Tibshirani, 1996), forward stepwise regression (Hocking, 1976), or the graphical Lasso (Friedman et al., 2008), with a tuning parameter that controls the model complexity, and select the value of the tuning parameter by minimizing the cross-validation (CV) estimate of out-of-sample prediction error. This paradigm is versatile but prioritizes prediction error above all else, while ignoring the other goals mentioned above. In particular, CV can be quite liberal in including irrelevant explanatory variables (see e.g. Baumann, 2003), an undesirable feature in applications where accurate variable selection is a priority.
@@
In such applications, we could abandon estimation methods like the Lasso and instead apply multiple testing methods such as knockoffs (Barber and Candès, 2016; Candès et al., 2018), or the Benjamini–Hochberg procedure or its dependence-corrected variants (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001; Fithian and Lei, 2022), which guarantee control of the false discovery rate (FDR) at a user-specified level. However, multiple testing methods do not yield fitted models, and the model defined by including only the selected variables can have very poor fit to the data.
@@
In this work we introduce a new approach for assessing variable selection accuracy of estimation methods like the Lasso: an estimator for the FDR of any variable selection procedure, which can be evaluated alongside the cross-validation curve to inform the analyst about the tradeoff between variable selection accuracy and prediction accuracy over the entire regularization path. Our estimator can be applied in common statistical settings including the Gaussian linear model, the Gaussian graphical model, and the nonparametric model-X framework of Candès et al. (2018), requiring no additional statistical assumptions in these settings beyond those required for standard hypothesis tests. Its bias is non-negative in finite samples and typically small, and its standard error can be assessed using the bootstrap. In the special case of thresholding independent $p$-values for multiple hypothesis tests, our method is almost identical to the FDR estimation method advocated by Storey (2002), but slightly less conservative.
@@
Figure 1: Cross-validation MSE (blue), true FDR (black), and our FDR estimator (red), for Lasso regression in two scenarios with explanatory variables that are (a) independent, and (b) highly correlated. In scenario (a) the minimum-CV model (solid blue vertical line) has high FDR, while the one-standard-error rule (dashed vertical line) achieves a reasonably low FDR. In scenario (b) there is no model that simultaneously achieves good predictive performance and low FDR. In both scenarios, our FDR estimator successfully captures important information about variable selection performance that is not evident from the CV curve.
Figure 1 illustrates why our method is useful by contrasting two Lasso regression scenarios with very different tradeoffs between prediction error and FDR. In both scenarios, we generate $n=600$ data points from a Gaussian linear model with $d=200$ standard Gaussian explanatory variables, of which only $d_{1}=20$ are true signal variables with the same moderate signal strength. The only difference is that the explanatory variables are independent in the first scenario, but correlated in the second (specifically, they follow $\text{AR}(0.8)$ correlation). This difference dramatically changes the variable selection performance: When the explanatory variables are independent, it is possible to select a good predictive model if we use the one-standard-error rule (but not if we minimize CV error). When the explanatory variables are correlated, however, all models with reasonably good predictive error have FDR exceeding $40\%$. While the CV curve gives no indication of the variable selection performance in either scenario, our estimator is accurate in both. The next section explains how we calculate it.
@@ 1.1 FDR estimation in the Gaussian linear model
Benjamini and Hochberg (1995) define the false discovery rate (FDR) as the expectation of the false discovery proportion (FDP), defined here as the fraction of selected variables that are actually irrelevant. If $\mathcal{H}_{0}$ is the index set of all noise variables and $R=|\mathcal{R}|$ is the total number of selections, we have
@@
using the convention that $/0=0$ 0; i.e., $\textnormal{FDP}=0$ when no variables are selected.
Note that we can linearly decompose the FDR as
@@
Using this decomposition, a promising estimator of the FDR is
@@
The definition of our estimator $\widehat{\textnormal{FDR}}$ in (5) is highly general and makes no assumptions about the form of the variable selection procedure $\mathcal{R}$. As a result, it applies to any selection procedure including the Lasso, forward stepwise regression, LARS (Efron et al., 2004), the elastic net (Zou and Hastie, 2005), and many others.
Our method requires no further assumptions beyond the standard linear model assumptions in (1) required to justify the marginal $t$-tests for each variable. Indeed, in the common case where the method $\mathcal{R}$ is a function of $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}$ and $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only (which is true of all methods referenced above), the only statistical assumption required to justify the inequality (6) is that the $t$-statistic follows a $t_{n-d}$ distribution under $H_{j}$, conditional on $\boldsymbol{S}_{j}$. This assumption may hold approximately even when the parametric assumptions of the linear model fail.
As we see next, our method generalizes far beyond the Gaussian linear model to other parametric and nonparametric contexts, provided that our modeling assumptions supply an appropriate sufficient statistic $\boldsymbol{S}_{j}$ for each null hypothesis $H_{j}$. Section 2 defines our method generically and discusses additional applications including the graphical Lasso (Friedman et al., 2008) and nonparametric model-X assumptions introduced by (Candès et al., 2018). Section 3 discusses theoretical analysis of the standard error in stylized settings, as well as our bootstrap scheme for estimating the standard error, which is used to generate the error bars in all figures, and Sections 4 and 5 illustrate our method’s behavior in real data and simulation examples, respectively. Section 6 discusses computational issues, and Section 7 concludes.
@@ 2.1 Notation and statistical assumptions
Throughout the paper, we will use upper-case letters to represent variables and lower-case letters to denote individual observations. For example, $y_{i}$ is the $i$th observation of variable $Y$ and $x_{i,j}$ is the $i$th observation of variable $X_{j}$. We use the upper-case bold-face letter to denote a vector of all $n$ observations, e.g. $\boldsymbol{X}_{j}=(x_{1,j},\ldots,x_{n,j})^{\mkern-1.5mu\mathsf{T}}$. We use $[d]$ to denote the set $\{1,\ldots,d\}$. For $B\subseteq[d]$, we denote as $\boldsymbol{X}_{B}$ the block of matrix $\boldsymbol{X}$ including only the variables in $B$, and we use $\boldsymbol{X}_{-j}$ as a shorthand for $\boldsymbol{X}_{[d]\setminus\{j\}}$ and $\boldsymbol{X}_{-B}$ for $\boldsymbol{X}_{[d]\setminus B}$.
We denote by $\boldsymbol{D}$ a generic random object that encodes the entire data set, whose distribution $P$ we will assume lies in some known family $\mathcal{P}$. If $\mathcal{P}$ is a parametric family, we generically use $\boldsymbol{\theta}$ to denote its parameters, and $P_{\boldsymbol{\theta}}$ to denote the corresponding distribution. In the Gaussian linear model in Section 1.1, irrelevance of explanatory variable $X_{j}$ is represented by the null hypothesis $H_{j}:\boldsymbol{\theta}_{j}=0$, but other problem settings will involve other null hypotheses.
@@ 2.2 FDR estimation
Step 1: Estimating the first factor:
Crucially, note that the factor $\mathbb{E}[\mathbf{1}\{j\in\mathcal{R}\}/R]$ is only relevant when $j\in\mathcal{H}_{0}$. As a result, we can assume when estimating it that $H_{j}$ holds. Since $\boldsymbol{S}_{j}$ is sufficient for the submodel $H_{j}$, we can do so by Rao-Blackwellizing the integrand (Rao, 1945; Blackwell, 1947), obtaining
@@
For $j\in\mathcal{H}_{0}$, $\widehat{\textnormal{FDR}}_{j}^{*}$ is unbiased by construction, and it is the UMVU estimator for the first factor of $\textnormal{FDR}_{j}$ whenever $\boldsymbol{S}_{j}$ is complete (Lehmann and Scheffé, 1950).
Step 2: Estimating the second factor:
We could obtain a conservative estimator by simply upper-bounding the indicator $\mathbf{1}\{j\in\mathcal{H}_{0}\}\leq 1$, resulting in the estimator $\widehat{\textnormal{FDR}}^{*}=\sum_{j}\widehat{\textnormal{FDR}}_{j}^{*}$. However this estimator’s bias, while still non-negative, could be quite large. In particular, if some of the signal variables are selected with high conditional probability given $\boldsymbol{S}_{j}$, their excess contributions may be large.
@@
or $\phi_{j}(\boldsymbol{D})=1$ if $\mathbb{E}_{H_{j}}[\psi_{j}|\boldsymbol{S}_{j}]=0$. Note we have $\mathbb{E}_{H_{j}}[\phi_{j}\mid\boldsymbol{S}_{j}]=1$ almost surely.
Commonly there is a standard test for $H_{j}$, whose $p$-value $p_{j}$ is uniform and independent of $\boldsymbol{S}_{j}$ under $H_{j}$. Then, we can take $\psi_{j}=\mathbf{1}\{p_{j}>\zeta\}$ for $\zeta\in(0,1)$, and
@@
If the signal is strong enough for the test to reject $H_{j}$ with high probability at level $\zeta$, then variable $j$’s contribution will almost always be zero. We adopt (9) as the canonical form of $\phi_{j}$, using $\zeta=0.1$ for all simulations.
@@
We can write the canonical version of $\widehat{\textnormal{FDR}}_{j}$ more concretely as
@@
These two factors are independent if $p_{j}\perp\!\!\!\perp\boldsymbol{S}_{j}$, as is commonly the case.
@@
For any selection procedure $\mathcal{R}$, and for any estimator $\psi_{j}:\;\boldsymbol{D}\to[0,1]$ of $\mathbf{1}\{j\in\mathcal{H}_{0}\}$, the estimator $\widehat{\textnormal{FDR}}$ defined in (710) has non-negative bias:
@@ 2.4 Example model assumptions
Assume we observe $n$ independent realizations of a $d$-variate Gaussian random vector $X$ ($n>d$), with mean zero and positive-definite covariance matrix $\boldsymbol{\Sigma}$. In the corresponding precision matrix $\boldsymbol{\Theta}=\boldsymbol{\Sigma}^{-1}$, the entry $\boldsymbol{\Theta}_{jk}$ is zero if and only if $X_{j}$ and $X_{k}$ are conditionally independent given the other entries. In other words, the sparsity pattern of the undirected dependence graph $\mathcal{G}$ for $X$ is identical to the sparsity pattern of $\boldsymbol{\Theta}$. In lieu of estimating all $d(d-1)/2$ off-diagonal entries of $\boldsymbol{\Theta}$, a common modeling assumption is that the dependence graph is sparse, and by extension that $\boldsymbol{\Theta}$ is sparse as well (Meinshausen and Bühlmann, 2006; Yuan and Lin, 2007; Friedman et al., 2008).

### v4

# Estimating the False Discovery Rate of Variable Selection

@@ 1 Introduction
When selecting variables for inclusion in a statistical model, data analysts typically seek to balance multiple goals, including predictive accuracy, reliable identification of signal variables, model parsimony, and interpretability. Perhaps the best-known paradigm is to use a variable selection method such as the Lasso (Tibshirani, 1996), forward stepwise regression (Hocking, 1976), or the graphical Lasso (Friedman et al., 2008),in conjunction with a tuning parameter that governs model complexity. The optimal value of this tuning parameter is commonly chosen by minimizing the cross-validation (CV) estimate of out-of-sample prediction error. Although this framework is flexible and effective, it inherently prioritizes prediction accuracy over other desirable properties. In particular, CV-based selection can be quite liberal in including irrelevant explanatory variables (see, e.g., Baumann, 2003), which is undesirable in applications where accurate variable selection is of primary importance.
@@
In such applications, one may forgo estimation-based approaches like the Lasso in favor of multiple testing procedures, such as knockoffs (Barber and Candès, 2016; Candès et al., 2018), or the Benjamini–Hochberg procedure and its dependence-corrected variants (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001; Fithian and Lei, 2022), which guarantee control of the false discovery rate (FDR) at a user-specified level. However, multiple testing methods do not yield fitted models, and the model formed by including only the selected variables may exhibit very poor fit to the data. [^4] For example, in the knockoff procedure, certain correlation structures can cause a strong signal variable to have a negative knockoff statistic with approximately 50% probability (Li and Fithian, 2021). If this occurs, the strong sigal variable will not be selected regardless of how large the FDR threshold is.
@@
In this work, we introduce a novel approach for assessing variable selection accuracy of estimation methods such as the Lasso: an estimator of the FDR applicable to any variable selection procedure under common statistical settings. This estimator can be evaluated in conjunction with the cross-validation curve, enabling analysts to examine the trade-off between variable selection accuracy and predictive performance across the entire regularization path. The proposed method is applicable in common statistical settings, including the Gaussian linear model, the Gaussian graphical model, and the nonparametric model-X framework of Candès et al. (2018), without imposing any additional assumptions beyond those required for standard hypothesis testing. The estimator exhibits non-negative finite-sample bias that is typically small, and its standard error can be assessed via the bootstrap.
The idea of estimating the FDR of a selection procedure dates back to Storey (2002), who proposed a conservative estimator for variable selection based on thresholding independent $p$-values. Goeman and Solari (2011) further developed an upper confidence bound for the number of null variables in any given selection set, yielding a conservative median estimator. Their method relies on a closed testing framework, and a feasible implementation typically requires independent or specific positively correlated $p$-values. In addition, certain multiple testing procedures, such as Benjamini–Hochberg (Benjamini and Hochberg, 1995) and knockoffs (Barber and Candès, 2016), incorporate intrinsic False Discovery Proportion (FDP) estimators to achieve FDR control based on their particular variable ordering and testing strategy. In contrast, our proposed FDR estimator is broadly applicable across common statistical settings and can be used with any variable selection procedure deemed appropriate by the analyst.
@@
Figure 1: Cross-validation MSE (blue), true FDR (black), and our FDR estimator (red), for Lasso regression in two scenarios with explanatory variables that are (a) independent, and (b) highly correlated. In scenario (a), the minimum-CV model (solid blue vertical line) has high FDR, while the one-standard-error rule (dashed vertical line) achieves a reasonably low FDR. In scenario (b), there is no model that simultaneously achieves good predictive performance and low FDR. In both scenarios, our FDR estimator successfully captures important information about variable selection performance that is not evident from the CV curve.
Figure 1 illustrates the utility of our proposed method by comparing two Lasso regression scenarios that exhibit markedly different trade-offs between prediction error and FDR. In both scenarios, we generate $n=600$ observations from a Gaussian linear model with $d=200$ standard Gaussian explanatory variables, among which only $d_{1}=20$ are signal variables of equal moderate strength. The sole difference between the two settings lies in the dependence structure of the explanatory variables: in the first scenario, the variables are independent, whereas in the second, they exhibit an autoregressive $\text{AR}(1)$ correlation structure with a correlation coefficient $0.8$. This distinction substantially alters the variable selection performance. When the explanatory variables are independent, a model exhibiting both high predictive accuracy and reliable variable selection can be obtained using the one-standard-error rule, though not by minimizing the CV error. In contrast, when the variables are correlated, all models achieving competitive predictive accuracy exhibit an FDR exceeding $40\%$. While the CV curve alone offers no insight into variable selection performance in either case, our proposed estimator remains accurate across both scenarios. The following section introduces its construction.
@@ 1.1 FDR estimation in the Gaussian linear model
Benjamini and Hochberg (1995) define the false discovery rate (FDR) as the expectation of the false discovery proportion (FDP), defined here as the fraction of selected variables that are actually irrelevant. Let $\mathcal{H}_{0}$ be the index set of all noise variables and $R=|\mathcal{R}|$ be the total number of selections, we have
@@
using the convention that $0/0=0$; i.e., $\textnormal{FDP}=0$ when no variables are selected.
Inspired by Fithian and Lei (2022), we linearly decompose the FDR as
@@
This technique allows us to treat each term separately and decouple the complicated interactions by conditioning. A promising estimator of the FDR is
@@
The definition of our estimator $\widehat{\textnormal{FDR}}$ in (5) is highly general and imposes no assumptions on the form of the variable selection procedure $\mathcal{R}$. Consequently, it is applicable to any selection method, including the Lasso, forward stepwise regression, LARS (Efron et al., 2004), the elastic net (Zou and Hastie, 2005), among many others.
Our approach requires no assumptions beyond those of the standard linear model in (1), which are required to justify the use of marginal $t$-tests for each variable. In the common setting where the procedure $\mathcal{R}$ depends only on $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}$ and $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$–a property shared by all methods listed above–the sole statistical assumption needed to justify the inequality (6) is that the $t$-statistic follows a $t_{n-d}$ distribution under $H_{j}$, conditional on $\boldsymbol{S}_{j}$. Notably, this assumption may still hold approximately even when the parametric assumptions of the linear model are violated.
As we show next, our method extends far beyond the Gaussian linear model to other parametric and nonparametric settings, provided that the modeling assumptions yield an appropriate sufficient statistic $\boldsymbol{S}_{j}$ for each null hypothesis $H_{j}$. Section 2 describes our method in full generality and discusses additional applications, including the graphical Lasso (Friedman et al., 2008) and the nonparametric model-X framework of Candès et al. (2018). Section 3 presents a theoretical analysis of the standard error in stylized settings and describes our bootstrap procedure for estimating the standard error, which is used to construct the error bars in all figures. Sections 4 and 5 illustrate the behavior of our estimator in real-data and simulation studies, respectively. Section 6 addresses computational considerations, and Section 7 concludes.
@@ 2.1 Notation and statistical assumptions
Throughout the paper, we use uppercase letters to represent random variables and lowercase letters to denote individual observations. For example, $y_{i}$ denotes the $i$th observation of the variable $Y$, and $x_{i,j}$ denotes the $i$th observation of the variable $X_{j}$. We use bold uppercase letters to denote the collection of all $n$ observations. For instance, $\boldsymbol{X}_{j}=(x_{1,j},\ldots,x_{n,j})^{\mkern-1.5mu\mathsf{T}}$ represents the vector of observations for the random variable $X_{j}$, and $\boldsymbol{X}\in\mathbb{R}^{n\times d}$, whose $(i,j)$th element is $x_{i,j}$, denotes the matrix of observations for all explanatory variables. Moreover, let $d]={1,\ldots,d}$. For any subset $A\subseteq[d]$, we denote by $\boldsymbol{X}_{A}$ the submatrix of $\boldsymbol{X}$ containing only the columns corresponding to the variables in $A$. Similarly, we use $\boldsymbol{X}_{-j}$ as shorthand for $\boldsymbol{X}_{[d]\setminus{j}}$ and $\boldsymbol{X}_{-A}$ for $\boldsymbol{X}_{[d]\setminus A}$.
We denote by $\boldsymbol{D}$ a generic random object representing the entire dataset, whose distribution $P$ is assumed to belong to a known family $\mathcal{P}$. If $\mathcal{P}$ is a parametric family, we use $\boldsymbol{\theta}$ to denote its parameters and write $P_{\boldsymbol{\theta}}$ for the corresponding distribution. In the Gaussian linear model discussed in Section 1.1, the irrelevance of explanatory variable $X_{j}$ is expressed through the null hypothesis $H_{j}:\boldsymbol{\theta}_{j}=0$, while other problem settings may involve different forms of null hypotheses. Moreover, we denote by $\mathcal{R}$ the index set of variables selected by a generic variable selection procedure based on the random dataset $\boldsymbol{D}$, and define $R=|\mathcal{R}|$ as the total number of selected variables.
@@ 2.2 FDR estimation
Step 1: Estimating the first factor. Crucially, note that the factor $\mathbb{E}[\mathbf{1}\{j\in\mathcal{R}\}/R]$ is only relevant when $j\in\mathcal{H}_{0}$. Therefore, when estimating it, we may assume that $H_{j}$ holds. Since $\boldsymbol{S}_{j}$ is sufficient for the submodel under $H_{j}$, we can estimate the first factor by Rao-Blackwellizing the integrand (Rao, 1945; Blackwell, 1947), obtaining
@@
For $j\in\mathcal{H}_{0}$, $\widehat{\textnormal{FDR}}_{j}^{*}$ is unbiased by construction, and it is the UMVU estimator for the first factor of $\textnormal{FDR}_{j}$ whenever $\boldsymbol{S}_{j}$ is complete (Lehmann and Scheffé, 1950). Since the conditional null distribution of the data $\boldsymbol{D}$ given $\boldsymbol{S}_{j}$ is known by the sufficiency, $\widehat{\textnormal{FDR}}^{*}_{j}$ is computable and can generally be approximated using a Monte Carlo algorithm. For certain variable selection procedures, we have developed fast and exact computational methods. Further details are provided in Section 6.
Step 2: Estimating the second factor. We could obtain a conservative estimator by simply upper-bounding the indicator $\mathbf{1}\{j\in\mathcal{H}_{0}\}\leq 1$, resulting in the estimator $\widehat{\textnormal{FDR}}^{*}=\sum_{j}\widehat{\textnormal{FDR}}_{j}^{*}$. However this estimator’s bias, while still non-negative, could be quite large. In particular, if some of the signal variables are selected with high conditional probability given $\boldsymbol{S}_{j}$, their excess contributions may be large.
@@
or $\phi_{j}(\boldsymbol{D})=1$ if $\mathbb{E}_{H_{j}}[\psi_{j}|\boldsymbol{S}_{j}]=0$. Note we have $\mathbb{E}_{H_{j}}[\phi_{j}\mid\boldsymbol{S}_{j}]=1$ almost surely. For the same reason as in Step 1, the normalizer $\mathbb{E}_{H_{j}}[\psi_{j}(\boldsymbol{D})\mid\boldsymbol{S}_{j}]$ is always computable.
As an example, commonly, there is a standard test for $H_{j}$, whose $p$-value $p_{j}$ is uniform and independent of $\boldsymbol{S}_{j}$ under $H_{j}$. Then, we can take $\psi_{j}=\mathbf{1}\{p_{j}>\zeta\}$ for $\zeta\in(0,1)$, and
@@
The parameter $\zeta$ governs the bias–variance trade-off: a larger $\zeta$ leads to a smaller bias but a larger variance. Our numerical experiments indicate that $\zeta=0.1$ provides a good balance between bias and variance. When the signal is strong enough for the test to reject $H_{j}$ with high probability at level $\zeta$, the contribution of variable $j$ will almost always be zero. In Appendix A.3, we establish an upper bound on the bias contributed by a strong-signal variable in the Gaussian linear model, showing that it decays exponentially with the signal strength. We adopt (9) as the canonical form of $\phi_{j}$.
@@
When we adopt the canonical $\phi_{j}$, we can write the canonical version of $\widehat{\textnormal{FDR}}_{j}$ more concretely as
@@
These two factors are independent when $p_{j}\perp\!\!\!\perp\boldsymbol{S}_{j}$. [^5] When a suitable $p_{j}$ that is uniform and independent of $\boldsymbol{S}_{j}$ under $H_{j}$ is not available to adopt the canonical form, researchers may instead choose an appropriate $\psi_{j}$ function and employ the general formulation (10) of $\widehat{\textnormal{FDR}}$.
@@
Suppose for each $j=1,\ldots,d$, $\boldsymbol{S}_{j}$ is a sufficient statistic under the null submodel $H_{j}$. Then for any selection procedure $\mathcal{R}$, and for any estimator $\psi_{j}:\;\boldsymbol{D}\to[0,1]$ of $\mathbf{1}\{j\in\mathcal{H}_{0}\}$, the estimator $\widehat{\textnormal{FDR}}$ defined in (7, 8, 10) has non-negative bias:
@@ 2.4 Example model assumptions
Assume we observe $n$ independent realizations of a $d$-variate Gaussian random vector $X$ ($n>d$), with mean zero and positive-definite covariance matrix $\boldsymbol{\Sigma}$. In the corresponding precision matrix $\boldsymbol{\Theta}=\boldsymbol{\Sigma}^{-1}$, the entry $\boldsymbol{\Theta}_{jk}$ is zero if and only if $X_{j}$ and $X_{k}$ are conditionally independent given the other entries. In other words, the sparsity pattern of the undirected dependence graph $\mathcal{G}$ for $X$ is identical to the sparsity pattern of $\boldsymbol{\Theta}$. In lieu of estimating all d(d1)/2d(d-1)/2 off-diagonal entries of $\boldsymbol{\Theta}$, a common modeling assumption is that the dependence graph is sparse, and by extension that $\boldsymbol{\Theta}$ is sparse as well (Meinshausen and Bühlmann, 2006; Yuan and Lin, 2007; Friedman et al., 2008).